# Bayesian Low-Rank LeArning (Bella): A Practical Approach to Bayesian Deep Learning

### v4

@@ Abstract
Bayesian learning has demonstrated merits in robustness and resilience to unseen and out-of-distribution inputs, yet its adoption in large-scale tasks remains limited. In particular, deep ensemble methods (Seligmann et al. 2024; Lakshminarayanan, Pritzel, and Blundell 2017) have proven to be highly effective, but they suffer from a high computational cost. This paper introduces an alternative framework that mitigates the computational burden of ensemble Bayesian deep learning. The approach is inspired by the recent success of low-rank adapters and is named Bayesian Low-Rank LeArning (Bella). We show that (1) Bella achieves a dramatic reduction in the number of trainable parameters required to approximate a Bayesian posterior; and (2) it not only maintains, but in some instances surpasses the performance—in accuracy and out-of-distribution generalisation—of conventional Bayesian and non-Bayesian methods. An extensive empirical evaluation on large-scale tasks (ImageNet, CAMELYON17, DomainNet, VQA) with CLIP and LLaVA demonstrates the effectiveness and versatility of Bella in building scalable and Bayesian deep models for real-world applications.
@@ 1 Introduction
Bayesian deep learning (Neal 2012) provides mechanisms for building predictive models that are more robust to adversarial attacks, resilient to unseen and out-of-distribution inputs, and a principled approach for estimating model uncertainty (Liu et al. 2019; Ye and Zhu 2018; Wilson et al. 2022; Izmailov et al. 2020; Chen and Ghattas 2020). In particular, quantifying uncertainty enables more reliable decision-making and facilitates identifying potential model vulnerabilities in uncertain regions of the input space. Therefore, Bayesian deep learning (BDL) has a high potential in helping build more reliable and trustworthy AI systems for various real-world applications (e.g., autonomous driving, medical image analysis, etc.). Unfortunately, BDL methods suffer from a high computational complexity.
Bayesian models contrast with traditional point estimates, i.e. a single set of parameters representing a single mapping from inputs to outputs. A Bayesian model, instead, learns a distribution of model parameters that captures a distribution over possible predictions. Let us consider a neural network $f(\mathbf{x},\boldsymbol{\theta})$ with input $\mathbf{x}$ parameterized by $\boldsymbol{\theta}$ and its corresponding prior distribution $p(\boldsymbol{\theta})$. The likelihood $p(\mathcal{D}|\boldsymbol{\theta})$ is determined by $f$ (Bishop 2006) and Bayesian inference uses Bayes’ theorem to derive a posterior distribution $p\left(\boldsymbol{\theta}|\mathcal{D}\right)={p\left(\mathcal{D}|\boldsymbol{\theta}\right)p\left(\boldsymbol{\theta}\right)}/{p(\mathcal{D})}$ and then compute the predictive distribution:
@@
Figure 1: Training (left) and test (right) error landscapes on the CAMELYON17 dataset (%). The left figure shows a learned 3-particle approximation of the modes of the posterior of a pre-trained model $\boldsymbol{\theta}$. The test landscape is slightly different from the training one, highlighting the challenge of generalization. While a point estimate does not generalize well (a single parameter particle such as $\boldsymbol{\theta}+\Delta\boldsymbol{\theta}_{1}$) Bella shows a clear benefit. The Bayesian prediction (Equation 1) effectively averages over multiple parameter settings and leads to lower test error.
Despite evidence for the benefits of Bayesian learning, its adoption is hindered by several challenges, most importantly the intractability of the posterior distribution $p(\boldsymbol{\theta}|\mathcal{D})$ in most practical tasks. The exact solution is impractical to compute even for networks of moderate size, hence the necessity of using approximations.
This paper expands upon recent research demonstrating the effectiveness of ensemble methods and Stein Variational Gradient Descent (SVGD) (Liu and Wang 2016) approaches in BDL (Seligmann et al. 2024; Lakshminarayanan, Pritzel, and Blundell 2017; Abbasnejad et al. 2020a; Doan et al. 2022; Izmailov et al. 2020; Chen and Ghattas 2020). These methods aim to provide a practical method for ensemble Bayesian deep learning for large-scale tasks. With the goal of achieving a practical and efficient approach, we consider spawning parameter particles (models) by adding linear interpolation to a pre-trained model’s parameters to an ensemble. To further enhance the efficiency of these interpolations, we take inspiration from low-rank adaptation (LoRAs)(Hu et al. 2022) and propose to use low-rank parameters in the interpolation process. We therefore name the approach Bayesian Low-Rank Learning (Bella).
We validated Bella by comparing it to the full-parameter alternative in terms of (1) accuracy, (2) out-of-distribution generalization, (3) adversarial robustness, and (4) alignment of uncertainty estimation with human confidence. In summary, our findings indicate that our approach achieves comparable performance to full SVGD or ensemble methods at a fraction of the computational cost, with the exception of adversarial robustness. For example, Figure 1 illustrates the application of our proposed approach of Bayesian inference in an application with distribution shifts (CAMELYON17 (Koh et al. 2021) benchmark). The particles (i.e., parameter samples from the posterior) obtained using Bella prove effective in improving generalization, improving the common single particle fine-tuning and comparable with full ensemble-based baselines at a fraction of the cost. Our key contributions are:
@@
  • We demonstrate Bella to perform on par or better than baselines—ensembles or current SVGD implementations. Bella consistently outperforms the non-BNN counterparts on 8 datasets; image classification (ImageNet, CIFAR 10/100), Out-of-distribution (DomainNet, Camleyon17, CIFAR-10-C), and VQA as well as adversarial robustness. Notably, it sets a new state-of-the-art on CAMELYON17 (see Table 7 and the Camelyon leaderboard (2024)).
  • Bella employed with the multi-modal model LlaVa (Liu et al. 2023b) leads to improved performance and uncertainty estimation highly correlated with human confidence.
@@ 2 Related Work
Parameter-Efficient Fine Tuning. In contrast to fine-tuning all parameters, recent research proposed inserting adapters in between existing neural layers to reduce the number of trainable parameters and, subsequently, the compute (GPU consumption) (Houlsby et al. 2019; Rebuffi, Bilen, and Vedaldi 2017; Lin, Madotto, and Fung 2020). Hu et al. (Hu et al. 2022) use a bottleneck structure to impose a low-rank constraint on the weight updates, named LoRA. The key functional difference is that LoRA can be merged with the main weights during inference, thus avoiding the introduction of any latency whilst significantly reducing the number of parameters.
@@
Fine-Tuning Approaches and Bayesian Deep Learning.  Previous research investigating the application of fine-tuning approaches for BDLs have predominantly focused on large language models (LLMs), e.g.  (Fan et al. 2020; Zhang et al. 2021; Yang et al. 2024). Notably, marking a departure from the conventional methods relying primarily on tuning the network’s parameters, these studies chose to define priors and approximate posterior over low rank attention weights. Concurrently, Yang et al. (Yang et al. 2024) introduced the concept of Laplace LoRA to incorporate Bayesian concepts to enhance the calibration of fine-tuned LLMs. However, Laplace’s method (Daxberger et al. 2021) relies on a Gaussian approximation of the posterior distribution. Whilst this can be effective for unimodal and symmetric distributions, the approach does not fully encapsulate the intricacies of more complex posteriors, particularly in neural networks where the posterior has multimodality and asymmetry (Izmailov et al. 2021). Deep ensembles (Lakshminarayanan, Pritzel, and Blundell 2017) typically perform better in practice compared with variational and Laplace methods, due to their ability to capture multiple modes. When employing fine-tuning, a direct application of ensembling for LoRAs was considered in (Wang, Aitchison, and Rudolph 2024). Some interpretations of deep ensembles suggest that they approximate gradient flows in function spaces and that building desirable properties into an ensemble (such as repulsive behavior), is possible (Wild et al. 2023). SVGD (Liu and Wang 2016), can be viewed in a similar vein. However, while these more sophisticated, repulsive, ensembling approaches are highly impractical in the pre-training phase, we argue that their expressivity can be brought to bear precisely in tandem with low-rank fine-tuning, which is the viewpoint we adopt in this contribution.
@@
Building upon these foundations, our research represents a pioneering effort to apply the principles of repulsive ensemble-based low-rank fine-tuning to computer vision. In particular, we bridge pre-training and fine-tuning phases with recent conjectures on mode connectivity (Ainsworth, Hayase, and Srinivasa 2023). Our methodology not only capitalizes on the efficiency of fine-tuning techniques, e.g. (Ding et al. 2023, 2022; Dettmers et al. 2023) but also innovatively addresses the scalability challenges inherent to BDLs. Overall, our work sets a new precedent in applying Bayesian approaches to computer vision tasks by offering a scalable and efficient framework for enhancing model performance.
3 Background on
Stein Variational Gradient Descent (SVGD)
@@ 3 Background on
Bayesian inference techniques have been integral to the development of neural networks, with a rich history underscored by previous works (Neal et al. 2011; Neal 2012; MacKay 1991; Welling and Teh 2011; Blei, Kucukelbir, and McAuliffe 2017). Variational Inference (VI) (Blundell et al. 2015; Blei, Kucukelbir, and McAuliffe 2017) and Markov Chain Monte Carlo (MCMC) (Neal et al. 2011; Welling and Teh 2011) are two primary approximate Bayesian inference frameworks. The former substitutes the true posterior with a tractable alternative while the latter involves sampling. However, accurately computing the posterior with either MCMC or VI becomes computationally infeasible when dealing with large-scale networks containing millions of parameters. Although approximations can be obtained more efficiently with VI (Blundell et al. 2015), VI is also demonstrably too restrictive to resemble the multi-modality of the true posterior and suffers from mode collapse (Izmailov et al. 2021).
@@
Stein Variational Gradient Descent (SVGD) (Liu and Wang 2016), among other ensemble-based approaches, is an alternative approximate Bayesian technique that combines the strengths of MCMC and VI by transporting a set of parameter particles to fit the true posterior distribution, while encouraging diversity among the particles, by incorporating a repulsive term in the parameter updates. This diversity prevents the mode collapse and enables learning multiple models to represent various patterns in the data. Using $n$ samples from the posterior (i.e. parameter particles), SVGD modifies the gradient descent as:
@@
4 Bayesian Low-Rank Learning (Bella)
The problem with the current SVGD and other ensemble-based methods in large deep neural networks is its huge computational cost. This renders it infeasible to train efficiently and to scale to a sufficient number of parameter particles for accurately approximating the posterior distribution, which currently remains coarse. In this work, we propose to capitalize on the low-rank representations of fine-tuning in order to construct a practical and scalable variant of SVGD. We note that while our approach may not fully capture the diversity of the multi-modal posterior, a recent conjecture (Ainsworth, Hayase, and Srinivasa 2023) suggests that these modes might result from parameter permutations in neural networks, leading to models that are functionally equivalent. Building on this idea, Bella could serve as a practical alternative with sufficient theoretical justification for ensemble methods.
@@ 4 Bayesian Low-Rank Learning (Bella)
where $\mathbf{B}_{i}\in\mathbb{R}^{d_{1}\times r}$, $\mathbf{A}_{i}\in\mathbb{R}^{r\times d_{2}}$ are the low-dimensional update parameters, and $r\ll d_{1},d_{2}$ is the rank of the update. Now Bella proceeds as the joint SVGD on $(\mathbf{A}_{i},\mathbf{B}_{i})$, with updates
@@
where we denote $k_{i,j}=k\left(\mathbf{B}_{j}\mathbf{A}_{j},\mathbf{B}_{i}\mathbf{A}_{i}\right)$. Here, we have placed a zero-mean Gaussian prior on $(\mathbf{A}_{i},\mathbf{B}_{i})$, but other choices are possible. Note that the kernel function on $(\mathbf{A}_{i},\mathbf{B}_{i})$ is given by $k\left(\boldsymbol{\theta}_{0}+\mathbf{B}_{j}\mathbf{A}_{j},\boldsymbol{\theta}_{0}+\mathbf{B}_{i}\mathbf{A}_{i}\right)$, which ensures that the similarity is computed on the original parameter space, and in the commonly used case of shift-invariant kernels, this simplifies to $k\left(\mathbf{B}_{j}\mathbf{A}_{j},\mathbf{B}_{i}\mathbf{A}_{i}\right)$. Further simplifications are obtained for specific kernel functions – in particular, in the case of Gaussian RBF, while the naive implementation would require the cost of $O(rd_{1}d_{2})$ for a single kernel evaluation, we can bring it down to $O(r^{2}(d_{1}+d_{2}))$ using standard trace manipulation, as described in the Appendix G. This procedure can be repeated across all dense layers.
@@
5 Experiments and Results
In this section, we provide an in-depth overview of our experimental setup, detailing the methodology, equipment, and procedures employed in the implementation of Bella. We aim to compare Bella with established ensemble-based methods and SVGD, both of which have demonstrated effectiveness in BDLs (Seligmann et al. 2024; Lakshminarayanan, Pritzel, and Blundell 2017).
5.1 Experimental Setup
Datasets. In this research, we have employed a variety of datasets, each selected for their relevance and contribution, they include CIFAR-10, CIFAR-100 (Krizhevsky, Hinton et al. 2009), CIFAR-10-C (Hendrycks and Dietterich 2019), STL-10 (Coates, Ng, and Lee 2011), CAMELYON17 (Bandi et al. 2018), ImageNet (Russakovsky et al. 2015), and DomainNet (Peng et al. 2019). We also consider VQA v2 dataset utilized for Visual Question Answering (VQA). Additional details are in Appendix I and Appendix H.

### v5

@@ Abstract
Computational complexity of Bayesian learning is impeding its adoption in practical, large-scale tasks. Despite demonstrations of significant merits such as improved robustness and resilience to unseen or out-of-distribution inputs over their non-Bayesian counterparts, their practical use has faded to near insignificance. In this study, we introduce an innovative framework to mitigate the computational burden of Bayesian neural networks (BNNs). Our approach follows the principle of Bayesian techniques based on deep ensembles, but significantly reduces their cost via multiple low-rank perturbations of parameters arising from a pre-trained neural network. Both vanilla version of ensembles as well as more sophisticated schemes such as Bayesian learning with Stein Variational Gradient Descent (SVGD), previously deemed impractical for large models, can be seamlessly implemented within the proposed framework, called Bayesian Low-Rank LeArning (Bella). In a nutshell, i) Bella achieves a dramatic reduction in the number of trainable parameters required to approximate a Bayesian posterior; and ii) it not only maintains, but in some instances, surpasses the performance of conventional Bayesian learning methods and non-Bayesian baselines. Our results with large-scale tasks such as ImageNet, CAMELYON17, DomainNet, VQA with CLIP, LLaVA demonstrate the effectiveness and versatility of Bella in building highly scalable and practical Bayesian deep models for real-world applications.
Code: https://bnn-bella.github.io/BNN-Bella/
@@ 1 Introduction
Figure 1: Train and Test Error (%) landscapes of CAMELYON17. Training landscape demonstrates the learned 3 particle approximation of the modes of the posterior from a pre-trained model $\boldsymbol{\theta}$—modification of parameters in the constrained region ($\Delta\boldsymbol{\theta}$) leads to approaching posterior modes in inference. Here, we observe a key benefit of our Bella approximation compared to a point estimate—while a single parameter particle (e.g. $\boldsymbol{\theta}+\Delta\boldsymbol{\theta}_{1}$) do not generalize well, Bayesian prediction with Equation 1, effectively an average over multiple parameter settings, leads to better performance.
Bayesian deep learning [1] provides mechanisms, for building predictive models more robust to adversarial attacks, resilient to unseen or out-of-distribution data and a theoretical framework for estimating model uncertainty [2, 3, 4, 5, 6, 7, 8]. Consequently, embracing Bayesian deep learning (BDL) represents a significant stride towards building more reliable and trustworthy AI systems for various real-world applications (e.g., autonomous driving, medical image analysis, etc.). Unfortunately, their practical use is encumbered by their computational complexity.
Unlike traditional alternatives with point estimates—a single set of model parameters mapping inputs to outputs—Bayesian models learn the distribution of model parameters to offer a distribution over possible predictions. Consider a neural network $f(\mathbf{x},\boldsymbol{\theta})$ with input $\mathbf{x}$ parameterized by $\boldsymbol{\theta}$ and its corresponding prior distribution $p(\boldsymbol{\theta})$. The likelihood $p(\mathcal{D}|\boldsymbol{\theta})$ is determined by $f$ [9] and then, Bayesian inference seeks to use Bayes’ theorem to derive a posterior distribution given by $p\left(\boldsymbol{\theta}|\mathcal{D}\right)={p\left(\mathcal{D}|\boldsymbol{\theta}\right)p\left(\boldsymbol{\theta}\right)}/{p(\mathcal{D})}$ to compute the predictive distribution:
@@
Despite evidence to support the adoption of Bayesian learning, their widespread use is hindered by several challenges; a primary issue is the intractability of the posterior distribution $p(\boldsymbol{\theta}|\mathcal{D})$ in practical learning tasks. The exact solution for the posterior, even for networks of moderate size, is impractical. Because of deep neural network’s complexity and the high-dimensional integral of the resulting denominator. This necessitates the use of approximations.
Our work expands upon recent research demonstrating the effectiveness of ensemble methods and Stein Variational Gradient Descent (SVGD) [10] approaches in BDL, for example, [11, 12, 13, 14, 4, 5], to provide a practical method for ensemble Bayesian deep learning for large-scale tasks. To achieve a practical and efficient approach, we consider spawning parameter particles (models) by adding linear interpolation to a pre-trained model’s parameters to build an ensemble. To further enhance the efficiency of these interpolations, inspired by the success of low-rank adapters (LORAs)[15], we propose using low-rank parameters in the interpolation process, we refer to as Bayesian Low-Rank Learning (Bella).
To validate this, we compare Bella to the full-parameter alternative in terms of (1) accuracy, (2) out-of-distribution generalization, (3) adversarial robustness, and (4) alignment of uncertainty estimation with human confidence. In summary, our findings indicate that this approach achieves comparable performance to full SVGD or ensemble methods at a fraction of the computational cost, with the exception of adversarial robustness. For example, Figure 1 illustrates the application of our proposed approach of Bayesian inference in an application with distribution shifts (CAMELYON17 [16] benchmark). The particles (i.e., parameter samples from the posterior) obtained using Bella prove effective in improving generalization, improving the common single particle fine-tuning and comparable with full ensemble-based baselines at a fraction of the cost. Our key contributions are:
@@
  • We demonstrate Bella performs on par or better than baselines—ensembles or current SVGD implementations. Bella consistently outperforms the non-BNN counterparts on 8 datasets; image classification (ImageNet, CIFAR 10/100), Out-of-distribution (DomainNet, Camleyon17, CIFAR-10-C), and VQA as well as adversarial robustness. Notably, it sets a new state-of-the-art for the CAMELYON17 (see Appendix for more details) according to the leaderboard [17].
  • Bella employed with the multi-modal model LlaVa [18], leads to improved performance and uncertainty estimation highly correlated with human confidence.
@@ 2 Related Work
Parameter-Efficient Fine Tuning. In contrast to fine-tuning all parameters, recent research proposed inserting adapters in between existing neural layers to reduce the number of trainable parameters and, subsequently, the compute time (GPU consumption) [19, 20, 21]. Hu et al. [15] use a bottleneck structure to impose a low-rank constraint on the weight updates, named LoRA. The key functional difference is that LoRA can be merged with the main weights during inference, thus avoiding the introduction of any latency whilst significantly reducing the number of parameters.
@@
Fine-Tuning Approaches and Bayesian Deep Learning.  Previous research investigating the application of fine-tuning approaches for BDLs have predominantly focused on large language models (LLMs), e.g.  [22, 23, 24]. Notably, marking a departure from the conventional methods relying primarily on tuning the network’s parameters, these studies chose to define priors and approximate posterior over low rank attention weights. Concurrently, Yang et al. [24] introduced the concept of Laplace LoRA to incorporate Bayesian concepts to enhance the calibration of fine-tuned LLMs. However, Laplace’s method [25] relies on a Gaussian approximation of the posterior distribution. Whilst this can be effective for unimodal and symmetric distributions, similar to [26, 27, 28], the approach does not fully encapsulate the intricacies of more complex posteriors, particularly in neural networks where the posterior has multimodality and asymmetry [29]. Deep ensembles [12] typically perform better in practice compared with variational and Laplace methods, due to their ability to capture multiple modes. When employing fine-tuning, a direct application of ensembling for LoRAs was considered in [30]. Some interpretations of deep ensembles suggest that they approximate gradient flows in function spaces and that building desirable properties into an ensemble (such as repulsive behavior), is possible [31]. SVGD [10], can be viewed in a similar vein. However, while these more sophisticated, repulsive, ensembling approaches are highly impractical in the pre-training phase, we argue that their expressivity can be brought to bear precisely in tandem with low-rank fine-tuning, which is the viewpoint we adopt in this contribution.
@@
Building upon these foundations, our research represents a pioneering effort to apply the principles of repulsive ensemble-based low-rank fine-tuning to computer vision. In particular, we bridge pre-training and fine-tuning phases with recent conjectures on mode connectivity [32]. Our methodology not only capitalizes on the efficiency of fine-tuning techniques, e.g. [33, 34, 35] but also innovatively addresses the scalability challenges inherent to BDLs. Overall, our work sets a new precedent in applying Bayesian approaches to computer vision tasks by offering a scalable and efficient framework for enhancing model performance.
3 Stein Variational Gradient Descent Primer
@@
Bayesian inference techniques have been integral to the development of neural networks, with a rich history underscored by previous works [36, 1, 37, 38, 39]. Variational Inference (VI) [40, 39] and Markov Chain Monte Carlo (MCMC) [36, 38] are two primary approximate Bayesian inference frameworks. The former substitutes the true posterior with a tractable alternative while the latter involves sampling. However, accurately computing the posterior with either MCMC or VI becomes computationally infeasible when dealing with large-scale networks containing millions of parameters. Although approximations can be obtained more efficiently with VI [40], VI is also demonstrably too restrictive to resemble the multi-modality of the true posterior and suffers from mode collapse [29].
@@
Stein Variational Gradient Descent (SVGD) [10], among other ensemble-based approaches, is an alternative approximate Bayesian technique that combines the strengths of MCMC and VI by transporting a set of parameter particles to fit the true posterior distribution, while encouraging diversity among the particles, by incorporating a repulsive term in the parameter updates. This diversity prevents the mode collapse and enables learning multiple models to represent various patterns in the data. Using $n$ samples from the posterior (i.e. parameter particles), SVGD modifies the gradient descent as:
@@
4 Bayesian Low-Rank Learning (Bella) for SVGD
The problem with the current SVGD and other ensemble-based methods in large deep neural networks is its huge computational cost. This renders it infeasible to train efficiently and to scale to a sufficient number of parameter particles for accurately approximating the posterior distribution, which currently remains coarse. In this work, we propose to capitalize on the low-rank representations of fine-tuning in order to construct a practical and scalable variant of SVGD. We note that while our approach may not fully capture the diversity of the multi-modal posterior, a recent conjecture [32] suggests that these modes might result from parameter permutations in neural networks, leading to models that are functionally equivalent. Building on this idea, Bella could serve as a practical alternative with sufficient theoretical justification for ensemble methods.
@@ 4 Bayesian Low-Rank Learning (Bella) for SVGD
where $\mathbf{B}_{i}\in\mathbb{R}^{d_{1}\times r}$, $\mathbf{A}_{i}\in\mathbb{R}^{r\times d_{2}}$ are the low-dimensional update parameters, and $r\ll d_{1},d_{2}$ is the update’s rank. Now Bella proceeds as the joint SVGD on $(\mathbf{A}_{i},\mathbf{B}_{i})$, with updates
@@
where we denote $k_{i,j}=k\left(\mathbf{B}_{j}\mathbf{A}_{j},\mathbf{B}_{i}\mathbf{A}_{i}\right)$. Here, we have placed a zero-mean Gaussian prior on $(\mathbf{A}_{i},\mathbf{B}_{i})$, but other choices are possible. Note that the kernel function on $(\mathbf{A}_{i},\mathbf{B}_{i})$ is given by $k\left(\boldsymbol{\theta}_{0}+\mathbf{B}_{j}\mathbf{A}_{j},\boldsymbol{\theta}_{0}+\mathbf{B}_{i}\mathbf{A}_{i}\right)$, which ensures that the similarity is computed on the original parameter space, and in the commonly used case of shift-invariant kernels, this simplifies to $k\left(\mathbf{B}_{j}\mathbf{A}_{j},\mathbf{B}_{i}\mathbf{A}_{i}\right)$. Further simplifications are obtained for specific kernel functions – in particular, in the case of Gaussian RBF, while the naive implementation would require the cost of $O(rd_{1}d_{2})$ for a single kernel evaluation, we can bring it down to $O(r^{2}(d_{1}+d_{2}))$ using standard trace manipulation, as described in the Appendix. This procedure can be repeated across all dense layers.
@@
5 Empirical Experiments and Results
In this section, we provide an in-depth overview of our experimental setup, detailing the methodology, equipment, and procedures employed in the implementation of Bella. We aim to compare Bella with established ensemble-based methods and SVGD, both of which have demonstrated effectiveness in BDLs [11, 12].
                         Models Bseline Models (SVGD) Bella Models (SVGD) Single
$n=3$ $n=5$ $n=20$ $n=40$ $n=3$ $n=5$ $n=20$ $n=100$ $n=1$
Storage Consumption (MB) 1321 $2$ 8868 17735 436 439 460 572 433
Table 1: Computational cost to train different models based on CLIP architecture for different datasets. Notably, with SVGD Baseline Models, we can only train up to $n$=40 particles on a A6000 48 GB GPU, while we can increase to more than 100 parameter particles with our Bella method with negligible increase of GPU consumption. The grey columns correspond to costs (parameters) of models in Table 2.
5.1 Experimental Set-up
Datasets. In this research, we have employed a variety of datasets, each selected for their relevance and contribution, they include CIFAR-10, CIFAR-100 [41], CIFAR-10-C [42], STL-10 [43], CAMELYON17 [44], ImageNet [45], and DomainNet [46]. We also consider VQA v2 dataset utilized for Visual Question Answering (VQA). Additional details are in the Appendix.